Silver metal reacts with sulfur to form silver sulfide according to the following reaction: 2Ag(s) + S(s) → Ag2S(s) a. Identify the limiting reagent if 50.0 g Ag reacts with 10.0 g S. b. What is the theoretical yield in g of Ag2S produced from the reaction? c. What is the amount in g of the excess reactant expected to remain after the reaction? d. When the reaction occurred, the amount of Ag2S obtained was 45.0 g. What is the percent yield of the reaction? Step 1. Analyze the problem Step 2. Determine the limiting reagent. Step 3. Compute the theoretical yield. Step 4. Compute the percent yield



1
Expert's answer
2021-11-17T00:39:01-0500

2Ag(s) + S(s) → Ag2S(s)


For one mole of sulphur there should be two moles of silver for complete reaction.

Moles of silver = 50/108 = 0.46 moles

Moles of elemental sulphur = 10/32 = 0.3125 moles

For 0.3125 moles of sulphur there must be 0.6250mole of silver.

So silver is limiting Reagent.